OpenStudy (anonymous):

derivative of f(x)=sin^2(5x)/(3-4x)

OpenStudy (anonymous):

(10sin5xcos5x)(3-4x)+4sin^2(5x)/(16x^2-24x+9)

OpenStudy (anonymous):

\[f'(x)=\frac{(10\sin 5x \cos 5x)(3-4x)+4\sin^2(5x)}{16x^2-24x+9}\]